OpenStudy (anonymous):

the driver of a 1560kg hybrid automobile traveling 24m/s on a level paved road hits the brake to stop for a red light. determine the distanced needed to stop the car if the coefficient of friction between the car tires and road is 0.80. can someone tell me how to find the stopping distance applying newton's second law?

OpenStudy (anonymous):

stoping distance depends on the retardation of the vehiclr which in turn depends upon the coefficient of friction the frictional force causes a retardation of \[\mu g\] which is 0.8*10=8m/s^2 now use v=u+at here we need a final velocity = 0 0= 24-8t time t = 3s then use s=ut+1/2at^2 s= 24*3 -1/2*8*9 =36m

OpenStudy (anonymous):

okay thanks. you're substituting u for initial velocity correct?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

okay is the equation you gave me the same as the equation for newton's second law in component form?

OpenStudy (anonymous):

no they are just fundamental kinematic equations

OpenStudy (anonymous):

oh i see. how would apply newton's second law to this problem?

OpenStudy (anonymous):

how would i*

OpenStudy (anonymous):

I am not sure if I am right, but this is how I applied newton's second law to this problem. so when the car comes to a complete stop, that means the total net force = zero. So the force that's moving the car and the frictional force add up to zero. F that's moving the car + frictional force = 0 so F of the car = frictional force (umg) F = umg = 12,230.4 then with that force you can find the acceleration F / m = a (should be negative) then after finding the acceleration using Newton's second law, the kinematics can be used to find the time and the final velocity.

OpenStudy (anonymous):

that was the same thing i did except i skipped two steps